The sum of two natural numbers is 18. If the product of one number with the square of the

other is a maximum, find the numbers.


1
Expert's answer
2022-04-12T15:58:06-0400

ANSWER : one natural number is "6" the second is "12" .

EXPLANATION

Let one number be "n\in N" , then the second number is "(18-n)" .To determine for which "n" the product "n(18-n)^2" will be maximum consider the function "f(x)=x(18-x)^2" on the interval"[1,18]" . Since "f'(x)=(18-x)^2-2x(18-x)=(18-x)(18-x-2x)=3(18-x)(6-x)" and "(18-x)\geq 0" on the interval "[1,18]" , then "f'(x)>0" if "x \in [1,6)" and "f'(x)<0" if "x\in (6,18)" . Hence the function "f(x)" increases on the interval "[1,6]" and decreases on "[6,18]"

Thus, "f(1)<f(2)<...<f(6);" "f(6)>f(7)>...>f(18)" .

. So, "max _{[1,18]]}f(x) =f(6)" . Therefore , one natural number is "6" the second is "12" .